An empty bucket put under two faucets. If one faucet is turned on alone,the bucket fills in 6 minutes.if the other faucet is turned on alone,the bucket fills in 4 minutes.if both are turned on,how many seconds will it take to fill the bucket?
Answer by Edwin McCravy(20054)

An empty bucket put under two faucets. If one faucet is turned on alone,the bucket fills in 6 minutes.if the other faucet is turned on alone,the bucket fills in 4 minutes.if both are turned on,how many seconds will it take to fill the bucket?
Two ways, with and without algebra:

Without algebra:

In 12 minutes the first faucet can fill 2 buckets while the second one
fills 3 buckets.  That's 5 buckets in 12 minutes or 1 bucket in 12÷5 
or 2.4 minutes.

With algebra:

Let x be the answer
Make this chart:

               buckets filled     time    rate in buckets/minute
1st faucet           1             6
2nd faucet           1             4
both open            1             x 

Fill in the rates by dividing buckets by minutes:

               buckets filled     time    rate in buckets/minute
1st faucet           1             6             1/6
2nd faucet           1             4             1/4
both faucets         1             x             1/x


   Equation:

      1st faucet's rate + 2nd faucet's rate = their combined rate:

                    1/6 + 1/4 = 1/x

Multiply through by 12x

                      2x + 3x = 12
                           5x = 12
                            x = 2.4 minutes.

It's easier without algebre, but I assume you are taking algebra.

Edwin